A new soft drink is being market tested. A sample of 400 individuals participated in the taste test and 80 indicated they like the taste. Conduct the following hypothesis test at 5% significance level.

Null Hypothesis: At least 22% of the population will like the new soft drink.
Alternative Hypothesis: Less than 22% of the population will like the new soft drink.

Group of answer choices

Reject the null

None of these

Accept the null

Do not reject the null
Answer #1

Solution :

This is the left tailed test .

The null and alternative hypothesis is

H0 : p = 0.22

Ha : p < 0.22

= x / n = 80 / 400 = 0.2

Test statistic = z

= - P0 / [P0 * (1 - P0 ) / n]

= 0.2 - 0.22/ [(0.2 * 0.8) / 400]

= -0.966

P-value = 0.1670

= 0.05

P-value >

Do not reject the null